OpenStudy (anonymous):

Where am I going wrong here? 6/x - 5 = x 6 - 5x = x^2 0 = x2 + 5x -6 -5/2 +- sqrt((5/2)^2 + 6) -5/2 +- sqrt(49) -5/2 +- 7 -5/2 +- 14/2 wrong... answers should be x: 1, x: -6

OpenStudy (anonymous):

Is the problem 6/(x-5) or (6/x)-5?

OpenStudy (baldymcgee6):

You need to move the -5 over first, and then make it into a quadratic and factor

OpenStudy (baldymcgee6):

6/x = x+5 6= x^2+5 x^2 +5 -6 = 0 (x-1)(x+6)=0

OpenStudy (anonymous):

yep ^

OpenStudy (anonymous):

How did you find -1 and 6 though...?

OpenStudy (baldymcgee6):

By factoring the quadratic

OpenStudy (baldymcgee6):

(x-1)(x+6)=0 x = 1, -6

OpenStudy (anonymous):

I don't see how you get this though: (x-1)(x+6)=0

OpenStudy (baldymcgee6):

I'm sure deybow is typing up a long factoring explanation for you :)

OpenStudy (anonymous):

See that's what I tried to do with the (overly complicated?) formula, but I keep getting the wrong answer.

OpenStudy (baldymcgee6):

If you want you can use the quadratic formula, but there are faster factoring methods

OpenStudy (anonymous):

This is the formula I tried to use: x^2 + px +q = 0 x = -p/2 +- sqrt((p/2)^2 -q)

OpenStudy (baldymcgee6):

\[(-b \pm \sqrt(b^2 - 4ac))/2a\]

OpenStudy (anonymous):

Basically at this point you have x^2+5x-6 and you have to factor it. When you do factor it, you get (x+6)(x-1). Since it's quadratic, or since it has an x^2 in it, you know that there will be two separate parts in the answer that will each have an x in them, so you can start out by writing (x+-_)(x+-_) because the two x's would multiply together to make the x^2. Hopefully this makes sense so far. And the reason why it's + or - is because you have to figure out later which side has to be added or subtracted. Now the _'s in each set of parenthesis have to be numbers that will multiply to get -6 and add together to get 5. These numbers happen to be 6 and -1. So, you would put (x+6)(x-1). Now remember, this is all equal to zero, so the solutions will then be x=-6 and 1 because either of those values would make the equation come out to zero. Hopefully that makes sense :P

OpenStudy (baldymcgee6):

Told you he was typing up a nice long explanation for you :)

OpenStudy (anonymous):

haha it's hard to explain actually :\

OpenStudy (anonymous):

it is, but it's a very ambitious answer! very much appreciated

OpenStudy (anonymous):

ha thanks.

OpenStudy (anonymous):

I'll try your formula instead: (−b±(√b2−4ac))/2a I still *think* the one I used should work, but, idk.

OpenStudy (anonymous):

thanks, both of you!

OpenStudy (anonymous):

Yeah the quadratic formula is the longer method, and also the way you have to calculate the answer for any quadractics that can't be factored the easy way.

OpenStudy (anonymous):

But if you have any questions about the shortcut method I can try to explain more :P

OpenStudy (anonymous):

I need to learn the formulas first, I'll get to shortcuts when I've got this down. ;) thanks though